OpenStudy (anonymous):

a + b = 750 b is 66 more than a. Calculate the values of a and b. Show your method please :D

OpenStudy (anonymous):

HUHH IM 11 MAKE ME UNDERSTAND!! XX

OpenStudy (hoblos):

a+a+66=750 2a=684 a=342 b=a+66=408

OpenStudy (hoblos):

** b=a+66 a=b-66

OpenStudy (anonymous):

actually, hoblos is right

OpenStudy (anonymous):

im confusedd!!! xx :$

OpenStudy (anonymous):

ok a+66=b

OpenStudy (anonymous):

You can use a+66 in place of b in a+b=750

OpenStudy (anonymous):

so it would look like a+(a+66)=750

OpenStudy (anonymous):

can you solve for a ?

OpenStudy (anonymous):

noo... im USELESS!!!!!

OpenStudy (anonymous):

you need to isolate a

OpenStudy (anonymous):

subtract both sides by 66

OpenStudy (anonymous):

so you'll have a+a+66-66=750-66

OpenStudy (anonymous):

byeeeeeeeeeeee its too confusinggg

OpenStudy (anonymous):

bye